# Joint Transmit and Jamming Power Optimization for Secrecy in Energy Harvesting Networks: A Reinforcement Learning Approach

### v1

@@ Abstract
In this paper, we address the problem of joint allocation of transmit and jamming power at the source and destination, respectively, to enhance the long-term cumulative secrecy performance of an energy-harvesting wireless communication system until it stops functioning in the presence of an eavesdropper. The source and destination have energy-harvesting devices with limited battery capacities. The destination also has a full-duplex transceiver to transmit jamming signals for secrecy. We frame the problem as an infinite-horizon Markov decision process (MDP) problem and propose a reinforcement learning-based optimal joint power allocation (OJPA) algorithm that employs a policy iteration (PI) algorithm. Since the optimal algorithm is computationally expensive, we develop a low-complexity sub-optimal joint power allocation (SJPA) algorithm, namely, reduced state joint power allocation (RSJPA). Two other SJPA algorithms, the greedy algorithm (GA) and the naive algorithm (NA), are implemented as benchmarks. In addition, the OJPA algorithm outperforms the individual power allocation (IPA) algorithms termed individual transmit power allocation (ITPA) and individual jamming power allocation (IJPA), where the transmit and jamming powers, respectively, are optimized individually. The results show that the OJPA algorithm is also more energy efficient. Simulation results show that the OJPA algorithm significantly improves the secrecy performance compared to all SJPA algorithms. The proposed RSJPA algorithm achieves nearly optimal performance with significantly less computational complexity marking it the balanced choice between the complexity and the performance. We find that the computational time for the RSJPA algorithm is around 75 percent less than the OJPA algorithm.
@@ I Introduction
To maximize the long-term cumulative performance in EH wireless networks, the application of RL is considered in [13, 14, 5, 15]. In [13], a wireless sensor network is considered where the sensor node is equipped with EH capability and a finite data buffer. Optimal energy allocation for sensing and transmission for the sensor node is obtained by maximizing the total throughput over a finite horizon of time. The solution is achieved by formulating a finite-horizon MDP problem using the backward induction algorithm. In [14], a point-to-point wireless system with EH source equipped with an infinite energy queue is considered. An MDP problem is formulated to decide whether the EH source transmit or not in a given time slot. The objective is to maximize the average number of successfully delivered packets per time slot by the source. In [5], an EH transmitter with a finite battery capacity is considered. The MDP-based problem is formulated that maximizes the expected total transmitted data over the lifetime of the transmitter under the finite battery capacity constraint. The system in [15] considers a transmitter with finite data buffer and energy consumption in data sensing in addition to the system features assumed in [14, 5]. A joint energy allocation strategy for transmission and sensing is obtained to maximize the expected total amount of data transmitted until the transmitter stops functioning. Problems formulated in [14, 5, 15] are infinite-horizon MDP, whereas the solutions are provided using value iteration in [14, 15] or policy iteration (PI) in [5]. It is to be noted that the PI-based algorithms converge faster than the value iteration-based algorithms [16].
@@
To address the problem of secrecy through PLS, recently, a few works [21, 22, 23] have employed RL framework for solution. A large wireless network with multiple access points (APs), users, and eavesdroppers is considered in [21]. If an AP has no associated user to receive data, it works as a jammer. The joint optimal user association and power allocation to the APs are considered by maximizing the sum secrecy capacity of the users. The soft actor-critic algorithm from the RL framework is proposed as a solution. A reconfigurable intelligent surface (RIS)-aided wireless system with a base station (BS) and multiple users in the presence of multiple eavesdroppers is considered in [22]. To improve the secrecy rate of the system, a design problem is formulated to jointly optimize the BS’s beamforming and the RIS’s reflecting beamforming. A deep reinforcement learning (DRL)-based technique is proposed to achieve a secure beamforming policy in dynamic time-varying channel conditions. A multiple-input single-output (MISO) downlink system is considered in [23] where a BS with multiple transmit antennas communicates with multiple single-antenna devices with the help of an RIS. Legitimate devices are classified into trusted and untrusted devices, where the untrusted devices may potentially eavesdrop on the trusted devices. A deep deterministic policy gradient (DDPG)-based RL algorithm is proposed to obtain the joint optimal RIS phases and transmit beamforming by maximizing the sum secrecy rate of trusted devices while ensuring performance guarantee to all trusted and untrusted devices.
@@
None of the aforementioned works consider the EH capability at the nodes, which is essential for extending the lifespan of a network. To determine the potential impact of current decisions on the future secrecy performance due to energy limitation in EH networks, a communication system consisting of an EH source node, a full-duplex destination node, and a full-duplex active eavesdropper is considered in [24]. The destination node transmits an artificial noise while decoding the transmitted signal from the source only if the eavesdropper does not transmit a jamming signal. An optimal source transmit power decision policy is obtained to maximize the long-term secrecy rate using the value iteration algorithm in the RL framework.
@@
In the systems with EH nodes, where cumulative performance is optimized, secrecy was not a concern [14, 5, 15, 13]. Though [24] achieves secrecy through full-duplex destination jamming, the article only focuses on optimizing the transmit power of the source. In EH wireless networks where the full-duplex destination is jamming for secrecy, there is a necessity to optimize jamming power jointly with the source transmit power to enhance the long-term cumulative secrecy performance. If too little jamming power is assigned, the jammer cannot launch an effective jamming attack, conversely, excessive power to create jamming attacks can exhaust the jammer’s battery and increase self-interference. Both may lead to a decreased long-term cumulative secrecy performance. One should find the right balance by assigning the optimal amount of power to both the transmitter and the jammer jointly.
@@
  • We study, for the first time, the optimal joint power allocation (OJPA) problem for source transmission and destination jamming to maximize the long-term expected total transmitted secure bits in an EH wireless network until the network stops functioning using the RL framework. The problem is formulated as an infinite-horizon MDP as the lifetime of the proposed network is a random variable. The proposed OJPA algorithm utilizes the PI algorithm for the solution due to its faster convergence.
@@ II System Model
Consider an EH wireless communication system where a source node S is communicating with a destination node D in the presence of a passive eavesdropping node E, as illustrated in Fig. 1. We assume that both nodes S and D are equipped with an EH device, whereas node E is equipped with a regular power supply from the traditional power grid. Furthermore, node D operates in a full-duplex mode, and nodes S and E operate in a half-duplex mode. The full-duplex mode enables node D to receive data from node S and simultaneously perform jamming attacks on node E. Simultaneous jamming attack causes self-interference to signal reception at node D. We assume that a mechanism for self-interference cancellation (SIC) is in-place at node D; however, residual self-interference remains and is captured through a factor $0\leq\alpha\leq 1$, where $\alpha=1$ implies no SIC is performed and $\alpha=0$ indicates SI is canceled completely.
Transmission occurs in a time-slotted manner over the period of $K$ time slots (TSs), and each TS is indexed by $k\in\mathcal{K}=\{0,1,2,\ldots,K-1\}$. The TSs have identical duration of $T_{s}$ seconds. The transmission duration is $T_{x}\leq T_{s}$ seconds within each TS. Since EH nodes are susceptible to physical destruction or hardware failure, we assume the network lifetime $K$ is a random variable. Let $\Gamma\in[0,1)$ be the probability that the network remains operational throughout a given TS enduring physical damage or hardware malfunctions, where $\Gamma$ is constant for each of TS. Accordingly, the network lifetime $K$ can be modeled as a geometrically distributed random variable with mean $1/(1-\Gamma)$ [15, 5].
@@ II-A EH Model
We assume that the energy harvested by node S and node D in the $k$th TS are $H_{\textrm{S}}^{(k)}\in\mathcal{H}$ and $H_{\textrm{D}}^{(k)}\in\mathcal{H}$ energy units, respectively, where =$\mathcal{H}=\{0,E_{H}\}$ is the set of possible harvested energy units. We model the EH process in each TS at node S and node D as an independent and identically distributed Bernoulli process with probability $p$ and $q$, respectively [25]. Accordingly, in each TS $k\in\mathcal{K}$, the probability of harvesting energy of $E_{H}$ energy units at node S and node D is $\mathbb{P}[H_{\textrm{S}}^{(k)}=E_{H}]=p$ and $\mathbb{P}[H_{\textrm{D}}^{(k)}=E_{H}]=q$, respectively, and the probability of not harvesting any energy is $\mathbb{P}[H_{\textrm{S}}^{(k)}=0]=1-p$ and $\mathbb{P}[H_{\textrm{D}}^{(k)}=0]=1-q$, respectively.
@@
Note that the energy utilized for the signal transmission or the jamming attack during the $k$th TS can not exceed the amount of energy stored in corresponding batteries. Similarly, the storage of harvested energy is limited by the battery capacity. As a result, the energy levels $B_{\textrm{S}}^{(k+1)}$ and $B_{\textrm{D}}^{(k+1)}$ in the $(k+1)$-th TS is updated from the $k$th TS as
$B_{\textrm{S}}^{(k+1)}$ =$=\left\{\begin{array}[]{lcl}\min\{B_{\textrm{S}}^{(k)}-P_{\textrm{S}}^{(k)}T_{x}+E_{H},B_{\textrm{S}}^{\textrm{max}}\}\,\,\mbox{for}\,\,H_{\textrm{S}}^{(k)}=E_{H}\\ B_{\textrm{S}}^{(k)}-P_{\textrm{S}}^{(k)}T_{x}\hskip 82.51282pt\mbox{for}\,\,H_{\textrm{S}}^{(k)}=0\end{array}\right.$ (3)
$B_{\textrm{D}}^{(k+1)}$ =$=\left\{\begin{array}[]{lcl}\min\{B_{\textrm{D}}^{(k)}-P_{\textrm{D}}^{(k)}T_{x}+E_{H},B_{\textrm{D}}^{\textrm{max}}\}\,\,\mbox{for}\,\,H_{\textrm{D}}^{(k)}=E_{H}\\ B_{\textrm{D}}^{(k)}-P_{\textrm{D}}^{(k)}T_{x}\hskip 82.51282pt\mbox{for}\,\,H_{\textrm{D}}^{(k)}=0\end{array}\right.$ (6)
where $P_{\textrm{S}}^{(k)}\in\mathcal{P}$ and $P_{\textrm{D}}^{(k)}\in\mathcal{P}$ are the power transmitted by node S and node D for signal transmission and jamming attack, respectively, in the $k$th TS, $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{M}\}$ is the set of $M$ possible transmit power levels, and $P_{\textrm{S}}^{(k)}T_{x}\leq B_{\textrm{S}}^{(k)}$, $P_{\textrm{D}}^{(k)}T_{x}\leq B_{\textrm{D}}^{(k)}$.
@@ II-C Performance Metrics
$C_{\textrm{S}}=\mathbb{E}\left[\mathbb{E}_{K}\left[\sum_{k=0}^{K-1}C_{\textrm{S}}^{(k)}T_{x}\right]\right]\quad\text{bits},$ (11)
@@ III Problem Formulation
The objective of the considered EH wireless communication system is to maximize the expected total transmitted secure bits $C_{\textrm{S}}$ in (11) by optimally allocating $P_{\textrm{S}}^{(k)}$ and $P_{\textrm{D}}^{(k)}$ in each TS until the network stops functioning. The solution should consider the probability that the network remains operational at each TS, current battery energy level, EH rate, channel condition, and self-interference attenuation factor. Accordingly, we formulate the problem of finding the joint power allocation for transmitting and jamming power as
@@
$0\leq P_{\textrm{S}}^{(k)}\leq\frac{B_{\textrm{S}}^{(k)}}{T_{x}}$ (12c)
$0\leq P_{\textrm{D}}^{(k)}\leq\frac{B_{\textrm{D}}^{(k)}}{T_{x}}.$ (12d)
@@ IV-A Preliminaries
    =$=\left\{P_{\textrm{S}}^{(k)},P_{\textrm{D}}^{(k)}\mid 0\leq P_{\textrm{S}}^{(k)}\leq\frac{B_{\textrm{S}}^{(k)}}{T_{x}},0\leq P_{\textrm{D}}^{(k)}\leq\frac{B_{\textrm{D}}^{(k)}}{T_{x}}\right\}.$ (13)
@@
    where $\mathbb{P}[B_{\textrm{S}}^{(k+1)}\mid B_{\textrm{S}}^{(k)},H_{\textrm{S}}^{(k)},P_{\textrm{S}}^{(k)}]$ and $\mathbb{P}[B_{\textrm{D}}^{(k+1)}\mid B_{\textrm{D}}^{(k)},H_{\textrm{D}}^{(k)},P_{\textrm{D}}^{(k)}]$ are equal to 1 if (3) and (6) are satisfied, zero otherwise. We also have $\mathbb{P}[H_{\textrm{S}}^{(k)}]=p$ when $H_{\textrm{S}}^{(k)}=E_{H}$, $\mathbb{P}[H_{\textrm{S}}^{(k)}]=1-p$ when $H_{\textrm{S}}^{(k)}=0$, $\mathbb{P}[H_{\textrm{D}}^{(k)}]=q$ when $H_{\textrm{D}}^{(k)}=E_{H}$, and $\mathbb{P}[H_{\textrm{D}}^{(k)}]=1-q$ when $H_{\textrm{D}}^{(k)}=0$. If $\mathbb{P}[B_{\textrm{S}}^{(k+1)}\mid B_{\textrm{S}}^{(k)},H_{\textrm{S}}^{(k)},P_{\textrm{S}}^{(k)}]$ and $\mathbb{P}[B_{\textrm{D}}^{(k+1)}\mid B_{\textrm{D}}^{(k)},H_{\textrm{D}}^{(k)},P_{\textrm{D}}^{(k)}]$ are equal to zero, (14) also becomes zero indicating the impossibility of a transition from state $s^{(k)}$ to state $s^{(k+1)}$ while taking action $a^{(k)}$.

### v2

@@ Abstract
In this paper, we address the problem of joint allocation of transmit and jamming power at the source and destination, respectively, to enhance the long-term cumulative secrecy performance of an energy-harvesting wireless communication system until it stops functioning in the presence of an eavesdropper. The source and destination have energy-harvesting devices with limited battery capacities. The destination also has a full-duplex transceiver to transmit jamming signals for secrecy. We frame the problem as an infinite-horizon Markov decision process (MDP) problem and propose a reinforcement learning (RL)-based optimal joint power allocation (OJPA) algorithm that employs a policy iteration (PI) algorithm. Since the optimal algorithm is computationally expensive, we develop a low-complexity sub-optimal joint power allocation (SJPA) algorithm, namely, reduced state joint power allocation (RSJPA). Two other SJPA algorithms, the greedy algorithm (GA), and the naive algorithm (NA) are implemented as benchmarks. In addition, the OJPA algorithm outperforms the individual power allocation (IPA) algorithms termed individual transmit power allocation (ITPA) and individual jamming power allocation (IJPA), where the transmit and jamming powers, respectively, are optimized individually. The results show that the OJPA algorithm is also more energy efficient. Results also show that the OJPA algorithm significantly improves the secrecy performance compared to all SJPA algorithms. The OJPA algorithm also outperforms the secrecy performance of a genetic algorithm-based RL algorithm and a finite-horizon RL algorithm. The proposed RSJPA algorithm achieves nearly optimal performance with significantly less computational complexity marking it the balanced choice between the complexity and the performance. We find that the computational time for the RSJPA algorithm with considering only 505050 percent of the total number of states is around 757575 percent less than the OJPA algorithm.
@@ I Introduction
To maximize the long-term cumulative performance in EH wireless networks, the application of RL is considered in [13, 14, 5, 15]. In [13], a wireless sensor network is considered where the sensor node is equipped with EH capability and a finite data buffer. Optimal energy allocation for sensing and transmission for the sensor node is obtained by maximizing the total throughput over a finite horizon of time. The solution is achieved by formulating a finite-horizon Markov decision process (MDP) problem using the backward induction algorithm. In [14], a point-to-point wireless system with EH source equipped with an infinite energy queue is considered. An MDP problem is formulated to decide whether the EH source transmit or not in a given time slot. The objective is to maximize the average number of successfully delivered packets per time slot by the source. In [5], an EH transmitter with a finite battery capacity is considered. The MDP-based problem is formulated that maximizes the expected total transmitted data over the lifetime of the transmitter under the finite battery capacity constraint. The system in [15] considers a transmitter with a finite data buffer and energy consumption in data sensing in addition to the system features assumed in [14, 5]. A joint energy allocation strategy for transmission and sensing is obtained to maximize the expected total amount of data transmitted until the transmitter stops functioning. Problems formulated in [14, 5, 15] are infinite-horizon MDP, whereas the solutions are provided using value iteration in [14, 15] or policy iteration (PI) in [5]. It is to be noted that the PI-based algorithms converge faster than the value iteration-based algorithms [16].
@@
To address the problem of secrecy through PLS, recently, a few works [21, 22, 23, 24, 25, 26, 27] have employed RL framework for solution. A large wireless network with multiple access points (APs), users, and eavesdroppers is considered in [21]. If an AP has no associated user to receive data, it works as a jammer. The joint optimal user association and power allocation to the APs are considered by maximizing the sum secrecy capacity of the users. The soft actor-critic algorithm from the RL framework is proposed as a solution. A reconfigurable intelligent surface (RIS)-aided wireless system with a base station (BS) and multiple users in the presence of multiple eavesdroppers is considered in [22]. To improve the secrecy rate of the system, a design problem is formulated to jointly optimize the BS’s beamforming and the RIS’s reflecting beamforming. A deep reinforcement learning (DRL)-based technique is proposed to achieve a secure beamforming policy in dynamic time-varying channel conditions. A multiple-input single-output (MISO) downlink system is considered in [23] where a BS with multiple transmit antennas communicates with multiple single-antenna devices with the help of an RIS. Legitimate devices are classified into trusted and untrusted devices, where the untrusted devices may potentially eavesdrop on the trusted devices. A deep deterministic policy gradient (DDPG)-based RL algorithm is proposed to obtain the joint optimal RIS phases and transmit beamforming by maximizing the sum secrecy rate of trusted devices while ensuring performance guarantee to all trusted and untrusted devices.
In [24], a secure Visible Light Communication (VLC) system is considered, where multiple light fixtures serve as friendly jammers. A DRL algorithm is implemented to optimize the friendly jamming policy, assuming continuous state and action spaces. In [25], a smart cyber-attack scenario is examined, where attackers can dynamically select their attack methods, such as jamming or eavesdropping. An RL solution is employed to predict the attack strategies and intelligently determine whether artificial noise should be added to the transmitted signal. A Reconfigurable Intelligent Surface (RIS)-mounted Unmanned Aerial Vehicle (UAV)-assisted maritime communication system under jamming attacks is analyzed in [26]. To jointly optimize the transmission power of the base station, the placement of the UAV-RIS, and the RIS’s reflecting beamforming, a DRL-based approach is proposed. In [27], a UAV-aided Non-Orthogonal Multiple Access (NOMA) system is investigated for data collection from Transmission Devices (TDs) in the presence of an eavesdropping attack. A group of Auxiliary Devices (ADs) is deployed to provide cooperative jamming against the eavesdropper. A DRL-based online optimization algorithm is introduced to maximize the total secrecy capacity by jointly optimizing the power allocations of TDs and ADs.
@@
None of the aforementioned works that address secrecy of networks using RL approaches in [21, 22, 23, 24, 25, 26, 27] consider the EH capability at the source or the destination nodes, which is essential for extending the lifespan of a network. To determine the potential impact of current decisions on the future secrecy performance due to energy limitation in EH networks, a communication system consisting of an EH source node, a full-duplex destination node, and a full-duplex active eavesdropper is considered in [28]. A self-interference attenuation factor is considered at the full-duplex nodes reflecting the difficulty of fully suppressing own transmit signal. The destination node transmits an artificial noise while decoding the transmitted signal from the source only if the eavesdropper does not transmit a jamming signal. An optimal source transmit power decision policy is obtained to maximize the long-term secrecy rate using the value iteration algorithm in the RL framework.
@@
In the systems with EH nodes, where cumulative performance is optimized, secrecy was not a concern [14, 5, 15, 13]. Though [28] achieves secrecy through full-duplex destination jamming, the article only focuses on optimizing the transmit power of the source. In EH wireless networks where both the source and the destination rely solely on energy harvesting and the full-duplex destination is jamming for secrecy, there is a necessity to optimize jamming power jointly with the source transmit power to enhance the long-term cumulative secrecy performance. If too little jamming power is assigned, the jammer cannot launch an effective jamming attack, conversely, excessive power to create jamming attacks can exhaust the jammer’s battery and increase self-interference. Both may lead to a decreased long-term cumulative secrecy performance. One should find the right balance by assigning the optimal amount of power to both the transmitter and the jammer jointly.
@@
  • We study, for the first time, the optimal joint power allocation (OJPA) problem for source transmission and destination jamming to maximize the long-term expected total transmitted secure bits where both the source and destination are energy harvesting in an EH wireless network until the network stops functioning using the RL framework. The problem is formulated as an infinite-horizon MDP as the lifetime of the proposed network is a random variable. The proposed OJPA algorithm utilizes the PI algorithm for the solution due to its faster convergence.
@@
  • Additionally, we compare the secrecy performance of the proposed RL algorithms (OJPA and RSJPA) with that of a genetic algorithm-based RL algorithm. The performance of the OJPA algorithm (infinite-horizon) is also compared with that of a finite-horizon RL algorithm. The results show that the OJPA algorithm outperforms both the genetic algorithm-based and finite-horizon RL algorithms.
@@ II System Model
Consider an EH wireless communication system where a source node S is communicating with a destination node D in the presence of a passive eavesdropping node E, as illustrated in Fig. 1. We assume that both nodes S and D are equipped with an EH device that contains a rechargeable battery with limited storage capacity, whereas node E is equipped with a regular power supply from the traditional power grid. Furthermore, node D operates in a full-duplex mode, and nodes S and E operate in a half-duplex mode. The full-duplex mode enables node D to receive data from node S and simultaneously perform jamming attacks on node E. Simultaneous jamming attack causes self-interference to signal reception at node D. We assume that a mechanism for self-interference cancellation (SIC) is in-place at node D; however, residual self-interference remains and is captured through a factor $0\leq\alpha\leq 1$, where $\alpha=1$ implies no SIC is performed and $\alpha=0$ indicates SI is canceled completely [28]. We assume that a central processing unit (CPU) exists in the network which has access to global channel state and battery state information, and performs power allocation decisions with the help of this information.
Transmission occurs in a time-slotted manner over the period of $K$ time slots (TSs), and each TS is indexed by $k\in\mathcal{K}=\{0,1,\ldots,K-1\}$. The TSs have an identical duration of $T_{s}$ seconds [13, 15, 5, 8]. We assume that S has a sufficient amount of data available for transmission in each TS. Since EH nodes are susceptible to physical destruction or hardware failure, we assume the network lifetime $K$ is a random variable. Let $\Gamma\in[0,1)$ be the probability that the network remains operational throughout a given TS enduring physical damage or hardware malfunctions, where $\Gamma$ is constant for each of TS. Accordingly, the network lifetime $K$ can be modeled as a geometrically distributed random variable with mean $1/(1-\Gamma)$ [15, 5].
@@ II-A EH Model
We assume that the energy harvested by node S and node D in the $k$th TS are $H_{\textrm{S}}^{(k)}\in\mathcal{H}_{\textrm{S}}$ and $H_{\textrm{D}}^{(k)}\in\mathcal{H}_{\textrm{D}}$ energy units, respectively, where $\mathcal{H}_{\textrm{S}}=\{0,E_{\textrm{S}}\}$ and $\mathcal{H}_{\textrm{D}}=\{0,E_{\textrm{D}}\}$ is the set of possible harvested energy units. We model the EH event in each TS at node S and node D as an independent and identically distributed Bernoulli process with probability $p$ and $q$, respectively, independent of data transmission process [29]. Accordingly, in each TS $k\in\mathcal{K}$, the probability of harvesting energy of $E_{\textrm{S}}$ and $E_{\textrm{D}}$ energy units at node S and node D, respectively, is $\mathbb{P}[H_{\textrm{S}}^{(k)}=E_{\textrm{S}}]=p$ and $\mathbb{P}[H_{\textrm{D}}^{(k)}=E_{\textrm{D}}]=q$, respectively, and the probability of not harvesting any energy is $\mathbb{P}[H_{\textrm{S}}^{(k)}=0]=1-p$ and $\mathbb{P}[H_{\textrm{D}}^{(k)}=0]=1-q$, respectively.
@@
[…]
$B_{\textrm{S}}^{(k+1)}\hskip-2.84544pt$ =$=\hskip-2.84544pt\left\{\hskip-5.69046pt\begin{array}[]{lcl}\min\{B_{\textrm{S}}^{(k)}-P_{\textrm{S}}^{(k)}T_{s}+E_{\textrm{S}},B_{\textrm{S}}^{\textrm{max}}\}\,\,\mbox{for}\,\,H_{\textrm{S}}^{(k)}\hskip-2.84544pt=\hskip-2.84544ptE_{\textrm{S}}\\ B_{\textrm{S}}^{(k)}-P_{\textrm{S}}^{(k)}T_{s}\hskip 82.51282pt\mbox{for}\,\,H_{\textrm{S}}^{(k)}=0\end{array}\right.$ (3)
$B_{\textrm{D}}^{(k+1)}\hskip-2.84544pt$ =$=\hskip-2.84544pt\left\{\hskip-5.69046pt\begin{array}[]{lcl}\min\{B_{\textrm{D}}^{(k)}-P_{\textrm{D}}^{(k)}T_{s}+E_{\textrm{D}},B_{\textrm{D}}^{\textrm{max}}\}\,\,\mbox{for}\,\,H_{\textrm{D}}^{(k)}\hskip-2.84544pt=\hskip-2.84544ptE_{\textrm{D}}\\ B_{\textrm{D}}^{(k)}-P_{\textrm{D}}^{(k)}T_{s}\hskip 82.51282pt\mbox{for}\,\,H_{\textrm{D}}^{(k)}=0\end{array}\right.$ (6)
[…]
@@ II-C Performance Metrics
$\mu=\mathbb{E}\left[\mathbb{E}_{K}\left[\sum_{k=0}^{K-1}C_{\textrm{S}}^{(k)}T_{s}\right]\right]\quad\text{bits},$ (11)
@@ III Problem Formulation
The objective of the considered EH wireless communication system is to maximize the expected total transmitted secure bits $\mu$ in (11) by optimally allocating $P_{\textrm{S}}^{(k)}$ and $P_{\textrm{D}}^{(k)}$ in each TS until the network stops functioning. The solution should consider the probability that the network remains operational at each TS, current battery energy level, EH rate, channel condition, and self-interference attenuation factor. Accordingly, we formulate the problem of finding the joint power allocation for transmitting and jamming power as
@@
$0\leq P_{\textrm{S}}^{(k)}\leq\frac{B_{\textrm{S}}^{(k)}}{T_{s}}$ (12c)
$0\leq P_{\textrm{D}}^{(k)}\leq\frac{B_{\textrm{D}}^{(k)}}{T_{s}}.$ (12d)
@@ IV-A Preliminaries
    =$=\left\{P_{\textrm{S}}^{(k)},P_{\textrm{D}}^{(k)}\mid 0\leq P_{\textrm{S}}^{(k)}\leq\frac{B_{\textrm{S}}^{(k)}}{T_{s}},0\leq P_{\textrm{D}}^{(k)}\leq\frac{B_{\textrm{D}}^{(k)}}{T_{s}}\right\}.$ (13)
@@
    where $\mathbb{P}[B_{\textrm{S}}^{(k+1)}\mid B_{\textrm{S}}^{(k)},H_{\textrm{S}}^{(k)},P_{\textrm{S}}^{(k)}]$ and $\mathbb{P}[B_{\textrm{D}}^{(k+1)}\mid B_{\textrm{D}}^{(k)},H_{\textrm{D}}^{(k)},P_{\textrm{D}}^{(k)}]$ are equal to 1 if (3) and (6) are satisfied, zero otherwise. We also have $\mathbb{P}[H_{\textrm{S}}^{(k)}]=p$ when $H_{\textrm{S}}^{(k)}=E_{S}$, $\mathbb{P}[H_{\textrm{S}}^{(k)}]=1-p$ when $H_{\textrm{S}}^{(k)}=0$, $\mathbb{P}[H_{\textrm{D}}^{(k)}]=q$ when $H_{\textrm{D}}^{(k)}=E_{D}$, and $\mathbb{P}[H_{\textrm{D}}^{(k)}]=1-q$ when $H_{\textrm{D}}^{(k)}=0$. If $\mathbb{P}[B_{\textrm{S}}^{(k+1)}\mid B_{\textrm{S}}^{(k)},H_{\textrm{S}}^{(k)},P_{\textrm{S}}^{(k)}]$ and $\mathbb{P}[B_{\textrm{D}}^{(k+1)}\mid B_{\textrm{D}}^{(k)},H_{\textrm{D}}^{(k)},P_{\textrm{D}}^{(k)}]$ are equal to zero, (14) also becomes zero indicating the impossibility of a transition from state $s^{(k)}$ to state $s^{(k+1)}$ while taking action $a^{(k)}$.